Given the equation (x - 2)2 + (y - 6)2 = 9 of a circle, identify the center and the radius.
Answer by mananth(16949)

The center (h,k) =(2,6) and radius = sqrt(9)=3
by comparing with the equation
(x-h)^2+(y-k)^2=r^2